# Reconciling Kaplan and Chinchilla Scaling Laws

### v1

@@ Abstract
Kaplan et al. [2020] (‘Kaplan’) and Hoffmann et al. [2022] (‘Chinchilla’) studied the scaling behavior of transformers trained on next-token language prediction. These studies produced different estimates for how the number of parameters ($N$) and training tokens ($D$) should be set to achieve the lowest possible loss for a given compute budget ($C$). Kaplan: $N_{\text{optimal}}\propto C^{0.73}$, Chinchilla: $N_{\text{optimal}}\propto C^{0.50}$. This note finds that much of this discrepancy can be attributed to Kaplan counting non-embedding rather than total parameters, combined with their analysis being performed at small scale. Simulating the Chinchilla study under these conditions produces biased scaling coefficients close to Kaplan’s. Hence, this note reaffirms Chinchilla’s scaling coefficients, by explaining the cause of Kaplan’s original overestimation. 111 Code for Section 2 available: https://github.com/TeaPearce/Reconciling_Kaplan_Chinchilla_Scaling_Laws
@@ 1 Introduction
Set up. Kaplan studied relationships in terms of non-embedding parameters ($N_{\setminus E}$) and non-embedding compute ($C_{\setminus E}$), excluding the linear layers embedding the vocabulary and position indices ($N_{E}$). By contrast, Chinchilla studied total parameters ($N_{T}$) and total compute ($C_{T}$). Define,
@@
Approach overview. Our approach uses information and data from the Chinchilla and Kaplan studies to estimate the scaling laws that would emerge if the Chinchilla relationship had been expressed in terms of $N_{\setminus E}$ $C_{\setminus E}$, and this had been done at the smaller model sizes used in Kaplan.
We will see that for large $N_{T}$, $N_{E}$ becomes a negligible portion of the model’s parameters and compute cost. Hence in the large parameter regime the two coefficients directly conflict with each other. At smaller values of $N_{T}$, $N_{E}$ is not negligible (this is the regime considered in Kaplan’s study – 768 to 1.5B parameters). We find that at the smaller end of this range, the relationship between $N^{*}_{\setminus E}$ $C_{\setminus E}$ is not in fact a power law. However, fitting a “local” power law at this small scale, produces a coefficient that is close to Kaplan’s, and hence reconciles these two results.
@@
  • Step 4. Simulate synthetic data from the loss model over the model sizes used Kaplan. Fit a local power law for $N^{*}_{\setminus E}$ in terms of $C_{\setminus E}$.
@@
2 Analysis
@@
$N_{T}=N_{\setminus E}+\gamma N_{\setminus E}^{1/3},$ (7)
for some constant $\gamma>0$. Aside from having several nice properties (strictly increasing and $\lim_{N_{T}\to\infty}N_{T}=N_{\setminus E}$  [^2] Proof. $N_{T}=N_{\setminus E}+\gamma N_{\setminus E}^{1/3}\implies N_{T}/N_{\setminus E}=1+\gamma N_{\setminus E}^{-2/3}$. Examining the r.h.s., $\lim_{N_{T}\to\infty}1+\gamma N_{\setminus E}^{-2/3}=1$, hence we conclude on the l.h.s $\lim_{N_{T}\to\infty}N_{\setminus E}/N_{T}=1$ or $N_{\setminus E}=N_{T}$. ), it can be motivated from both the Kaplan and Chinchilla study.
@@
This is the same form as Eq. 7 with $\gamma=(v+h)\left(\frac{A}{12}\right)^{1/3}$.
@@
Chinchilla perspective. We empirically fit a function $N_{T}=N_{\setminus E}+\gamma N_{\setminus E}^{\delta}$ (note the learnable exponent) to the Chinchilla model configurations listed in Table A9 of Hoffmann et al. [2022] for a range of $N_{T}$ (44M to 16B). We compute $N_{E}$ from Eq. 2, using the reported vocab size of 32,000, but ignore the context length 2,048 since Chinchilla used non-learnable position embeddings (though their inclusion effects coefficients only slightly).
Figure 1 shows the configurations and relationship from a model fitted with numpy’s polyfit, which produces coefficients, $\gamma=47491$ $\delta=0.34$. The exponent has come out close to 1/3, and an implied aspect ratio $A=39.2$ (inferred from $\gamma$). Hence, this further supports the form in Eq. 7.
@@
$N^{*}_{T}=\underset{N_{T}\;\text{s.t.}\;C_{T}=6N_{T}D}{\operatorname{argmin}}\text{Loss}(N_{T},C_{T}).$ (11)
@@
$\text{Loss}(N_{T},D)$ $=\frac{N_{c}}{N_{T}^{\alpha}}+\frac{D_{c}}{D^{\beta}}+E,$ (12)
$\text{Loss}(N_{T},C_{T})$ $=\frac{N_{c}}{N_{T}^{\alpha}}+\frac{D_{c}}{(C_{T}/6N_{T})^{\beta}}+E,$ (13)
@@
$\text{Loss}(N_{\setminus E},C_{\setminus E})$ $=\frac{N_{c}}{(N_{\setminus E}+\gamma N_{\setminus E}^{1/3})^{\alpha}}+\frac{D_{c}}{(C_{\setminus E}/6N_{\setminus E})^{\beta}}+E$ (15)
@@
$6N^{*}_{\setminus E}\left(N^{*}_{\setminus E}+\frac{\gamma}{3}{N^{*}_{\setminus E}}^{1/3}\right)^{-\frac{1}{\beta}}\left(N^{*}_{\setminus E}+\gamma{N^{*}_{\setminus E}}^{1/3}\right)^{\frac{1+\alpha}{\beta}}\left(\frac{\beta}{\alpha}\frac{D_{c}}{N_{c}}\right)^{\frac{1}{\beta}}=C_{\setminus E}.$ (16)
@@
$\frac{1}{g}\coloneqq\frac{d\log(C_{\setminus E})}{d\log(N^{*}_{\setminus E})}=1-\frac{1}{\beta}\frac{{N^{*}_{\setminus E}}^{2/3}+\frac{\gamma}{9}}{{N^{*}_{\setminus E}}^{2/3}+\frac{\gamma}{3}}+\frac{\alpha+1}{\beta}\frac{{N^{*}_{\setminus E}}^{2/3}+\frac{\gamma}{3}}{{N^{*}_{\setminus E}}^{2/3}+\gamma}.$ (17)
Figure 2 plots Eq. 16 and 17, using coefficients for $\alpha,\beta,N_{c},D_{c}$ from the Epoch AI specification (see later), and $\gamma=47491$. There are three phases.
@@
  • At large scale, $\lim_{N^{*}_{\setminus E}\to\infty}\frac{1}{g}=\frac{\alpha+\beta}{\beta}\implies N^{*}_{\setminus E}\mathchoice{\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\displaystyle\propto\cr\kern 2.0pt\cr\displaystyle\sim\cr\kern-2.0pt\cr}}}}{\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\textstyle\propto\cr\kern 2.0pt\cr\textstyle\sim\cr\kern-2.0pt\cr}}}}{\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\scriptstyle\propto\cr\kern 2.0pt\cr\scriptstyle\sim\cr\kern-2.0pt\cr}}}}{\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\scriptscriptstyle\propto\cr\kern 2.0pt\cr\scriptscriptstyle\sim\cr\kern-2.0pt\cr}}}}C_{{\setminus E}}^{\frac{\beta}{\alpha+\beta}}$ 444 Proof. As $N^{*}_{\setminus E}\to\infty$, we can ignore $\gamma$ terms and $1/g=1-(1/\beta)+(\alpha+1)/\beta=1+\alpha/\beta$. , as in the $N_{T}$ case in Eq. 14.
  • There is also a transition phase, where $g$ briefly increases. This happens in between the two limits, when $N_{\setminus E}^{2/3}$ is of the same order as $\gamma$. Indeed at exactly the point $N_{\setminus E}^{2/3}=\gamma$, we have $N_{T}=N_{\setminus E}+\gamma N_{\setminus E}^{1/3}=N_{T}=2N_{\setminus E}$, or a 50/50 split between embedding and non-embedding parameters. In Figure 2 we see this transition region occurs around this point; $N_{\setminus E}=\gamma^{3/2}=47491^{3/2}\approx 1\times 10^{7}$.
@@
Step 4. Simulate synthetic data from the loss model over the model sizes used Kaplan. Fit a local power law for $N^{*}_{\setminus E}$ in terms of $C_{\setminus E}$.
By reading $g$ off Figure 2, we could estimate a local power law and hence scaling coefficient for a given value of $N^{*}_{\setminus E}$. However, it’s not clear what $N^{*}_{\setminus E}$ point value is representative of the Kaplan study. We opt for a more faithful estimation procedure, generating synthetic training curves from Eq. 15 across the range of model sizes used in Kaplan, and fit coefficients using models falling on the compute efficient frontier. This will also verifies our analytic expression for $N^{*}_{\setminus E}$ $C_{\setminus E}$ in Eq. 16.
Figure 3 shows the synthetic training curves generated. We simulated 20 models with $N_{\setminus E}$ ranging from 790 parameters to 1.58B (Kaplan reports using model sizes “ranging in size from 768 to 1.5 billion non-embedding parameters”). For other constants in Eq. 15, there are two options. Whilst the original Chinchilla study reported finding constants of $N_{c}=406.4,Dc=410.7,\alpha=0.3392,\beta=0.2849,E=1.6934\implies N^{*}_{T}\propto C_{T}^{0.46}$ (‘Chinchilla specification’), Besiroglu et al. [2024] conducted a re-analysis, arguing that the following set of constants were more accurate $N_{c}=482.01,Dc=2085.43,\alpha=0.3478,\beta=0.3658,E=1.8172\implies N^{*}_{T}\propto C_{T}^{0.51}$ (‘Epoch AI specification’). Our figures adopt the Epoch AI specification and $\gamma=47491$, though we report final results for the Chinchilla specification also.
@@
3 Experiments
@@
Result 2. Table 1 shows that optimization scheme has a smaller impact on scaling coefficients than switching from $N_{T}$ to $N_{\setminus E}$. Using a single set of models with no annealing (scheme 2) produces the same coefficients as using the more computationally expensive scheme 4. Counter to Chinchilla’s comment that moving from Kaplan’s scheme 3 to scheme 4 would reduce the scaling coefficient, our experiment suggests the opposite is the case, increasing from 0.46 to 0.49. This might explain our slight overestimation of the scaling coefficients in Eq. 26 & 35.
4 Discussion
This note aimed to explain the difference between the Kaplan and Chinchilla scaling coefficients. We found two issues in Kaplan’s study that combined to bias their estimated scaling coefficient; their choice to count only non-embedding parameters, and studying smaller sized model sizes. This means there is curvature in the true relationship between $N^{*}_{\setminus E}$ $N_{T}$ (Figure 4). At larger values of $N_{T}$, the embedding parameter counts become negligible, $N_{T}=N_{\setminus E}$, and differences would not arise. Alternatively, had Kaplan studied relationships directly in terms of $N_{T}$, this issue would also not arise, even at this smaller scale (confirmed by our Experiment 1 finding $N_{T}\propto C_{T}^{0.49}$ even for $N_{T}<5M$).
Why might embedding parameters be expected to contribute to scaling behavior? Although we do not have a definitive answer, several works evidence that embedding parameters capture meaningful language properties. Word embedding dimensions can be factorized into semantically interpretable factors [Mikolov et al., 2013a, b, Arora et al., 2018], and LLMs learn linear embeddings of space and time across scales [Gurnee and Tegmark, 2024]. Developing such meaningful embedding structures allows LLMs to perform high-level language operations, such as arithmetic [McLeish et al., 2024]. Therefore, if one believes that the embedding layer does more than just "translate" tokens to a vector of the correct dimension, we see no reason to exclude them in the parameter count. A direction for future work is to investigate the relative importance of each parameter type (MLP, QKV projections, embeddings etc.), and whether assigning a count of one to each type is correct.
Limitations. We acknowledge several limitations of our analysis. We have aimed to capture the primary ‘first order’ reason for the difference between the Kaplan and Chinchilla scaling coefficients. But there are multiple other differences between the two studies that likely also affect scaling coefficients; datasets (Kaplan used OpenWebText2, Chinchilla used MassiveText), transformer details (Kaplan used learnable position embeddings while Chinchilla’s were fixed, also differing tokenizers, vocabularly sizes), optimization scheme (Kaplan used scheme 3, Chinchilla scheme 4), differences in computation counting (Kaplan used $C=6ND$, Chinchilla’s Method 1 & 2 used a full calculation). However, our preliminary work suggested these factors impact coefficients in a more minor way.

### v2

@@ Abstract
Kaplan et al. [2020] (‘Kaplan’) and Hoffmann et al. [2022] (‘Chinchilla’) studied the scaling behavior of transformers trained on next-token language prediction. These studies produced different estimates for how the number of parameters ($N$) and training tokens ($D$) should be set to achieve the lowest possible loss for a given compute budget ($C$). Kaplan: $N_{\text{optimal}}\propto C^{0.73}$, Chinchilla: $N_{\text{optimal}}\propto C^{0.50}$. This paper finds that much of this discrepancy can be attributed to Kaplan counting non-embedding rather than total parameters, combined with their analysis being performed at small scale. Simulating the Chinchilla study under these conditions produces biased scaling coefficients close to Kaplan’s. Hence, this paper reaffirms Chinchilla’s scaling coefficients, by explaining the primary cause of Kaplan’s original overestimation. As a second contribution, the paper explains differences in the reported relationships between loss and compute. These findings lead us to recommend that future scaling studies use total parameters and compute.  [^1] Code for analysis: https://github.com/TeaPearce/Reconciling_Kaplan_Chinchilla_Scaling_Laws
Figure 1: Overview of the approach used to reconcile the two studies.
@@ 1 Introduction
We find that this methodological difference also plays a part in differences in the reported relationship between compute and loss. Kaplan: $L_{\text{optimal}}\propto C^{0.057}$, Chinchilla: $L_{\text{optimal}}-E\propto C^{0.178}$.
Concretely, this paper offers the following contributions.
  • We develop an analytical approach comparing the scaling relationships reported in Kaplan and Chinchilla.
  • This approach finds that Kaplan’s reported relationship is locally consistent with Chinchilla’s, if non-embedding parameters are used, and at smaller scale.
  • Secondly, we study the reported relationships between compute and loss. Again non-embedding parameters and smaller scale models are the cause of Kaplan’s biased estimate, combined with the absence of an offset term in their compute-loss form.
  • We recommend that going forward the scaling community measures total parameters and compute, and uses an offset in the compute-loss relationship.
2 Set Up
Kaplan studied relationships in terms of non-embedding parameters ($N_{\setminus E}$) and non-embedding compute ($C_{\setminus E}$), excluding the linear layers embedding the vocabulary and position indices ($N_{E}$). By contrast, Chinchilla studied total parameters ($N_{T}$) and total compute ($C_{T}$). Define,
@@ 2 Set Up
Our approach uses information and data from the Chinchilla and Kaplan studies to estimate the scaling laws that would emerge if the Chinchilla relationship had been expressed in terms of $N_{\setminus E}$ $C_{\setminus E}$, and this had been done at the smaller model sizes used in Kaplan, as summarized in Figure 1.
We will see that for large $N_{T}$, $N_{E}$ becomes a negligible portion of the model’s parameters and compute cost. Hence in the large parameter regime the two coefficients directly conflict with each other. At smaller values of $N_{T}$, $N_{E}$ is not negligible (this is the regime considered in Kaplan’s study – 768 to 1.5B parameters). We find that at the smaller end of this range, the relationship between $N^{*}_{\setminus E}$ $C_{\setminus E}$ is not in fact a power law. However, fitting a “local” power law at this small scale, produces a coefficient that is close to Kaplan’s, and hence roughly reconciles these two results.
@@
  • Step 4. Simulate synthetic data from the Chinchilla loss model over the model sizes used Kaplan. Fit a local power law for $N^{*}_{\setminus E}$ in terms of $C_{\setminus E}$.
@@
Section 5 presents a second, related contribution. We reconcile differences in the relationship between loss and compute proposed by the two studies. We leverage a similar analysis as above, but Step 3 & 4 are now performed with respect to the relationship between the optimal loss $L^{*}_{\setminus E}$, and compute $C_{\setminus E}$. So again, we begin from Chinchilla data, and correct for the exclusion of embedding parameters and compute, combined with the smaller model sizes used in Kaplan’s study, and additionally a differing choice of fitting function. Through these corrections, we are able to roughly recover Kaplan’s compute-loss coefficient, and hence reconcile the two studies.
3 Analysis: Compute-Parameter Scaling Coefficient
@@
$N_{T}=N_{\setminus E}+\omega N_{\setminus E}^{1/3},$ (7)
for some constant $\omega>0$. Aside from having several nice properties (strictly increasing and $\lim_{N_{T}\to\infty}N_{T}=N_{\setminus E}$  [^2] Proof. $N_{T}=N_{\setminus E}+\omega N_{\setminus E}^{1/3}\implies N_{T}/N_{\setminus E}=1+\omega N_{\setminus E}^{-2/3}$. Examining the r.h.s., $\lim_{N_{T}\to\infty}1+\omega N_{\setminus E}^{-2/3}=1$, hence we conclude on the l.h.s $\lim_{N_{T}\to\infty}N_{\setminus E}/N_{T}=1$ or $N_{\setminus E}=N_{T}$. ), it can be motivated from both the Kaplan and Chinchilla study.
@@
This is the same form as Eq. 7 with $\omega=(v+h)\left(\frac{A}{12}\right)^{1/3}$.
@@
Chinchilla perspective. We empirically fit a function $N_{T}=N_{\setminus E}+\omega N_{\setminus E}^{\delta}$ (note the learnable exponent) to the Chinchilla model configurations listed in Table A9 of Hoffmann et al. [2022] for a range of $N_{T}$ (44M to 16B). We compute $N_{E}$ from Eq. 2, using the reported vocab size of 32,000, but ignore the context length 2,048 since Chinchilla used non-learnable position embeddings (though their inclusion effects coefficients only slightly).
Figure 2 shows the configurations and relationship from a model fitted with numpy’s polyfit, which produces coefficients, $\omega=47491$ $\delta=0.34$. The exponent has come out close to 1/3, and an implied aspect ratio $A=39.2$ (inferred from $\omega$). Hence, this further supports the form in Eq. 7.
@@
$N^{*}_{T}=\underset{N_{T}\;\text{s.t.}\;C_{T}=6N_{T}D}{\operatorname{argmin}}L(N_{T},C_{T}).$ (11)
@@
$L(N_{T},D)$ $=\frac{N_{c}}{N_{T}^{\alpha}}+\frac{D_{c}}{D^{\beta}}+E,$ (12)
$L(N_{T},C_{T})$ $=\frac{N_{c}}{N_{T}^{\alpha}}+\frac{D_{c}}{(C_{T}/6N_{T})^{\beta}}+E,$ (13)
@@
$L(N_{\setminus E},C_{\setminus E})$ $=\frac{N_{c}}{(N_{\setminus E}+\omega N_{\setminus E}^{1/3})^{\alpha}}+\frac{D_{c}}{(C_{\setminus E}/6N_{\setminus E})^{\beta}}+E$ (15)
@@
$6N^{*}_{\setminus E}\left(N^{*}_{\setminus E}+\frac{\omega}{3}{N^{*}_{\setminus E}}^{1/3}\right)^{-\frac{1}{\beta}}\left(N^{*}_{\setminus E}+\omega{N^{*}_{\setminus E}}^{1/3}\right)^{\frac{1+\alpha}{\beta}}\left(\frac{\beta}{\alpha}\frac{D_{c}}{N_{c}}\right)^{\frac{1}{\beta}}=C_{\setminus E}.$ (16)
@@
$\frac{1}{g}\coloneqq\frac{d\log(C_{\setminus E})}{d\log(N^{*}_{\setminus E})}=1-\frac{1}{\beta}\frac{{N^{*}_{\setminus E}}^{2/3}+\frac{\omega}{9}}{{N^{*}_{\setminus E}}^{2/3}+\frac{\omega}{3}}+\frac{\alpha+1}{\beta}\frac{{N^{*}_{\setminus E}}^{2/3}+\frac{\omega}{3}}{{N^{*}_{\setminus E}}^{2/3}+\omega}.$ (17)
Figure 3 plots Eq. 16 and 17, using coefficients for $\alpha,\beta,N_{c},D_{c}$ from the Epoch AI specification (see later), and $\omega=47491$. There are three phases.
@@
  • At large scale, $\lim_{N^{*}_{\setminus E}\to\infty}\frac{1}{g}=\frac{\alpha+\beta}{\beta}\implies N^{*}_{\setminus E}\mathchoice{\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\displaystyle\propto\cr\kern 2.0pt\cr\displaystyle\sim\cr\kern-2.0pt\cr}}}}{\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\textstyle\propto\cr\kern 2.0pt\cr\textstyle\sim\cr\kern-2.0pt\cr}}}}{\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\scriptstyle\propto\cr\kern 2.0pt\cr\scriptstyle\sim\cr\kern-2.0pt\cr}}}}{\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\scriptscriptstyle\propto\cr\kern 2.0pt\cr\scriptscriptstyle\sim\cr\kern-2.0pt\cr}}}}C_{{\setminus E}}^{\frac{\beta}{\alpha+\beta}}$ 444 Proof. As $N^{*}_{\setminus E}\to\infty$, we can ignore $\omega$ terms and $1/g=1-(1/\beta)+(\alpha+1)/\beta=1+\alpha/\beta$. , as in the $N_{T}$ case in Eq. 14.
  • There is also a transition phase, where $g$ briefly increases. This happens in between the two limits, when $N_{\setminus E}^{2/3}$ is of the same order as $\omega$. Indeed at exactly the point $N_{\setminus E}^{2/3}=\omega$, we have $N_{T}=N_{\setminus E}+\omega N_{\setminus E}^{1/3}=N_{T}=2N_{\setminus E}$, or a 50/50 split between embedding and non-embedding parameters. In Figure 3 we see this transition region occurs around this point; $N_{\setminus E}=\omega^{3/2}=47491^{3/2}\approx 1\times 10^{7}$.
@@
Step 4. Simulate synthetic data from the Chinchilla loss model over the model sizes used Kaplan. Fit a local power law for $N^{*}_{\setminus E}$ in terms of $C_{\setminus E}$.
By reading $g$ off Figure 3, we could estimate a local power law and hence scaling coefficient for a given value of $N^{*}_{\setminus E}$. However, it’s not clear what $N^{*}_{\setminus E}$ point value is representative of the Kaplan study. We opt for a more faithful estimation procedure, generating synthetic training curves from Eq. 15 across the range of model sizes used in Kaplan, and fit coefficients using models falling on the compute efficient frontier. This will also verify our analytic expression for $N^{*}_{\setminus E}$ $C_{\setminus E}$ in Eq. 16.
Figure 4 shows the synthetic training curves generated. We simulated 20 models with $N_{\setminus E}$ ranging from 790 parameters to 1.58B (Kaplan reports using model sizes “ranging in size from 768 to 1.5 billion non-embedding parameters”). For other constants in Eq. 15, there are two options. Whilst the original Chinchilla study reported finding constants of $N_{c}=406.4,Dc=410.7,\alpha=0.3392,\beta=0.2849,E=1.6934\implies N^{*}_{T}\propto C_{T}^{0.46}$ (‘Chinchilla specification’), Besiroglu et al. [2024] conducted a re-analysis, arguing that the following set of constants were more accurate $N_{c}=482.01,Dc=2085.43,\alpha=0.3478,\beta=0.3658,E=1.8172\implies N^{*}_{T}\propto C_{T}^{0.51}$ (‘Epoch AI specification’). Our figures adopt the Epoch AI specification and $\omega=47491$, though we report final results for the Chinchilla specification also.
@@
4 Experiments: Compute-Parameter Scaling Coefficient
@@
Result 2. Table 1 shows that optimization scheme has a smaller impact on scaling coefficients than switching from $N_{T}$ to $N_{\setminus E}$. Using a single set of models with no annealing (scheme 2) produces the same coefficients as using the more computationally expensive scheme 4. Counter to Chinchilla’s comment that moving from Kaplan’s scheme 3 to scheme 4 would reduce the scaling coefficient, our experiment suggests the opposite is the case, increasing from 0.46 to 0.49. This might explain our slight overestimation of the scaling coefficients in Eq. 52 & 61.